OpenStudy (anonymous):

g(x)=sqrt(x) * f(x) find g'(4) given f(4)=3 and f'(4)=-5 i get -7/4 but my book says answer is -37/4

OpenStudy (anonymous):

\[g'(4)=\sqrt4\times f'(4)+f(4)\times \frac{1}{2\sqrt4}\]

OpenStudy (anonymous):

that be the product rule

OpenStudy (anonymous):

i know that but it does not equal the answer in my book